### v2

# ?reference?

Abstract
Automating the creation of scientific diagrams from academic papers can streamline the development of tutorials, presentations, and posters. Current text-to-image models struggle with generating accurate and visually appealing diagrams from long-context inputs. We propose SciDoc2Diagram, a task that extracts relevant information from scientific papers and generates diagrams, along with a benchmarking dataset, SciDoc2DiagramBench. We develop a multi-step pipeline SciDoc2Diagrammer that generates diagrams based on user intent using intermediate code generation. We observed that initial diagram drafts are often incomplete or unfaithful to the source, requiring SciDoc2Diagrammer-Multi-Aspect-Feedback (MAF), a refinement strategy that significantly enhances factual correctness and visual appeal and outperforms existing models on both automatic and human judgement.
1 Introduction
Researchers often need to transform detailed studies into compelling diagrams for presentations, going beyond the usual figures in their papers to include summaries, methodologies, and dynamic data visualizations Chapman et al. (2014). [^1] According to Fu et al. (2021), approximately 39.2% of figures in ACL slides are creatively derived from the original papers, taking forms like flowcharts, summary tables, and plots/charts that visually interpret the paper’s content. Automation of diagram generation from the multimodal components of academic papers—such as text, figures, tables, and plots—could enable the creation of accurate and appealing visual aids for scientific tutorials and presentations Mondal et al. (2024); Fu et al. (2021); Sun et al. (2021), which can help researchers and educators to improve productivity and enhancing the overall quality of their presentations. We introduce a new task SciDoc2Diagram where the input comprises of the content of academic papers and user’s intent and the output is a scientific diagram crafted from the content of papers and aligned with user’s intent, which we call “Extrapolated-Diagram”.
Figure 1: An example of Diagram Generation and Refinement using SciDoc2Diagrammer-MAF with input as a paper and user-defined intent.
Consider a scenario where a busy scientist wants to create figures from papers within a tight deadline, with the specific goal of generating a bar chart as shown in the Figure 1. To achieve this quickly, recent advancements in text-to-image (T2I) generation models like Stable Diffusion Rombach et al. (2022a) and DALL-E Ramesh et al. (2021, 2022a) offer promising solutions. However, these models face limitations in generating factually correct and visually appealing scientific figures. They primarily produce raster graphics at low resolutions, which is not suitable for scientific figures requiring precision and valid text Hsu et al. (2021). Belouadi et al. (2023) introduced a scientific diagram generation method based on only user’s intent. However, creation of informative diagrams using long-content papers as an additional input remains unexplored.
How effective are the current state-of-the-art Large Language models (LLMs) or Visual Language Models (VLMs) at generating scientific diagrams? Despite their potential, a significant knowledge gap exists: there is currently no standard benchmark to assess their performance in this area. To address this gap, we introduce the SciDoc2DiagramBench dataset that features 1080 diagrams from 89 *ACL papers, sourced from existing slides in ACL Anthology (Section 3). This dataset supports researchers in automating scientific diagram creation for presentations.
We evaluate state-of-the-art LLMs Jiang et al. (2023); Touvron et al. (2023); OpenAI (2023); Abdin et al. (2024) to extract relevant information from documents pertinent to the user-specified intent. Specifically, we use a multi-step pipelined method, SciDoc2Diagrammer that first extracts relevant data from academic papers corresponding to intent of the user in creating the diagram, generates an intermediate code that renders the final diagram (Section 4.1) as shown in Figure 1. However, the first-draft images generated by these models suffer from a range of errors: hallucination, incompleteness and unreadability due to their incapabilities of parsing long-prompts Lee et al. (2024a).
Therefore, we propose a sequential refinement approach SciDoc2Diagrammer-MAF based on the intial draft generated by SciDoc2Diagrammer. The new approach is guided by several critic models to provide targeted feedback (multiple aspects such as Completeness, Faithfulness, and Look and Feel) in a step-by-step hierarchical manner on the output to further enhance the quality of output (Section 4.2). Our multi-aspect feedback refinement approach significantly improves diagram quality, as confirmed by both automatic and human evaluations, demonstrating notable enhancements over existing methods (Section 6). Complex flowcharts and tables derived from single or multiple documents show notable enhancements in faithfulness and completeness when refined sequentially. It is particularly advantageous when iteratively synthesizing information from multiple sources to enhance clarity and precision. [^2] Code and data available at https://github.com/Ishani-Mondal/SciDoc2DiagramGeneration.
2 Task Formulation of SciDoc2Diagram
Our goal is to automatically generate scientific diagrams from academic papers tailored to user specifications. We process the input document $D=\{T,F,\tau,\pi\}$, which includes body text $T$, figures $F$, tables $\tau$, and plots $\pi$, guided by a user intent $I$ (e.g., “Create a bar chart showing the comparison…” in Figure 18). The output, an Extrapolated-Diagram $\delta$, visually represents content not explicitly illustrated in the paper.
Conversion Type # Diagrams
Extrapolated-Flowchart 320
Extrapolated-Results 290
Extrapolated-Architecture 270
Extrapolated-Summary 200
Total 1080
Table 1: Statistics of SciDoc2DiagramBench Dataset which includes 1080 diagrams and the distribution highlights the diversity of diagrams.
3 SciDoc2DiagramBench: A New Dataset
Creating Extrapolated-Diagram from source papers is an important yet challenging task. However, no current dataset evaluates state-of-the-art models on this task. To advance research, we have created SciDoc2DiagramBench dataset, repurposing DOC2PPT Fu et al. (2021) (scientific papers and corresponding slides from ACL Anthology) and TutorialBank Fabbri et al. (2018) (tutorials from *ACL conferences) [^3] Data Contamination issues are discussed in Appendix 14. Our dataset includes the following three parts and the overall statistics of different diagram types are in Table 1 with examples in Appendix 15.
SciDoc2DiagramBench-Gold.
Our initial goal was to explore what are the types of Extrapolated-Diagram usually created by humans while making presentations. We hired an expert annotator from Upwork with more than five years of authoring NLP-related papers and creating presentations. The annotator had two tasks, first to identify slides containing Extrapolated-Diagram (excludes slides that used external information not found in the paper), second to manually write an intent for the corresponding diagram creation (13, 12). Through this annotation process, we collect 265 gold-standard Extrapolated-Diagram from 34 papers where each instance comprises of a tuple of <papername, intent of diagram, gold diagram>. The first three authors manually checked the intents, and only instances that were unanimously deemed reasonable were retained. After analyzing these diagrams, 32% were flowcharts illustrating methodology (Extrapolated-Flowchart), 33% were results/plots (Extrapolated-Results), 25% were summary of related works/contributions (Extrapolated-Summary) and 10% were related to clarifying existing architecture (Extrapolated-Architecture).
SciDoc2DiagramBench-Extended.
Using insights from SciDoc2DiagramBench-Gold on how presenters transform paper content into diagrams, we used GPT-4 to generate new diagram intents for this subset of data (Table 14). We also provide guidelines for creating Extrapolated-Diagram that blend text and visuals from academic papers. Each intent, reviewed by the first author, aims to guide the creation of practical diagrams. Each instance includes <paperName, intent> but lacks a corresponding human-created gold diagram.
SciMultiDoc2DiagramBench-Gold.
To evaluate the task on another challenging set, we examined scientific ACL tutorials from TutorialBank Fabbri et al. (2018), focusing on slides where creators used content from multiple papers to design their diagrams. An expert annotator from Upwork identified these slides with Extrapolated-Diagrams. The first author then manually wrote intents for these tutorial slides, with each instance documented as <paperNames, intent, diagram>.
Figure 2: Outlines the procedure for generating diagrams from academic papers based on specific user intents, followed by refinement of each critic models.
4 Methodology
The SciDoc2Diagram task comprises of generating initial first-draft generation of diagrams from documents using base models (Section 4.1), followed by refinement of diagrams using critic modules (Section 4.2).
4.1 SciDoc2Diagrammer
We introduce SciDoc2Diagrammer, a two-stage pipeline for generating “Extrapolated-Diagrams” from paper content (Upper portion of Figure 2). In the first Diagram Planning stage, a planner VLM takes the PDF (decomposed into sections and figures/tables) and user’s intent and generates a diagram plan. Inspired by the success of extracting information using question-generation and retrieving answers from relevant snippets as done by Krithara et al. (2023); Lee et al. (2023); Qiu et al. (2018); Auer et al. (2023), we also synthesize diagram plan by first generating clarifying questions based on the user’s intent, and then obtain answers of those questions from relevant portions of the paper to come up with question-answer pairs, which we refer to as diagram plan. In the second stage, inspired by the success of LLMs in generating code from natural language instructions Ugare et al. (2024); Maddigan and Susnjak (2023), we use LLMs to take the diagram plan and user intent as inputs to generate an intermediate code and execute the code to finally render the diagram.
Diagram Planning:
Given an input document $D=\{T,F,\tau,\pi\}$ consisting of body text $T$, figures $F$, tables $\tau$, and plots $\pi$, and user intent $I$, the diagram planning phase includes the following sub-steps: [1] Intent Classification: We use an LLM (Few-shot with 4 exemplars) to classify if the intent of the user is related to results or not. For example in Figure 3, the intent is not related to results, but the one in Figure 2 is related to results. [2] Question Generator: Using the intent, we ask LLMs to generate clarification questions to create the diagram conditioned on intent. For example, for the intent in Figure 2, we obtain a set of clarification questions. [3] Retrieve Information: We retrieve relevant passages from the paper based on each of questions Lewis et al. (2020). In Figure 2, a plot and two paragraphs from Experimental Section of the source paper are retrieved to answer the generated questions. If any result type figure is relevant to the question, we first extract data from the figures and convert them into tables for simplification of content retrieval as motivated by prior works Liu et al. (2023a); Masry et al. (2023). [4] Diagram Plan Synthesis: After generating questions and extracting relevant content from paper, we extract answers of those questions. Then we formulate a dictionary of QA pairs/diagram plan $Q=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n})\}$, where $q_{i}$ are the questions and $a_{i}$ are the answers.
Diagram Rendering:
In this phase, we use the user intent $I$ and QA pairs $Q$ from the previous step to generate an intermediate code $\Theta$ that constructs diagram. This intermediate code is represented as Python code, as shown in Figure 2. Finally, we execute the intermediate code $\Theta$ to render “Extrapolated-diagram” $\delta$, as exemplified by the bar chart at the top-right corner of Figure 2.
Figure 3: The figure (an example from SciDoc2DiagramBench) depicts SciDoc2Diagrammer-MAF, which refines diagrams based on user intent and source document. Initially, SciDoc2Diagrammer creates a diagram (Step 2), which is refined through three critic modules which assess and provide feedback on necessary components, data accuracy, and visual design. The diagram is repeatedly refined, as shown on the left side of the figure (illustrating SumMAF), where feedback from all critics is integrated at Step 4’. The refinement continues, evaluated at Step 5’, until the maximum iterations are reached or the diagram meets the specified standards.
4.2 SciDoc2Diagrammer-MAF
Motivation: To gain a qualitative understanding of the diagrams generated from SciDoc2Diagrammer, we randomly sample 50 diagrams from the test set for manual inspection. There are some frequently occurring errors (Table 10) and we classify them into three broad categories: Completeness (essential components are missing or incomplete, affecting the diagram’s informativeness), Faithfulness (misleading information that do not accurately reflect the source data), and Layout (poor visual alignment hindering readability and comprehension). Besides, we also noticed that most of the errors propagate during the transition from our diagram planning to rendering stage; this was due to the inability of LLMs to handle long-context prompts Nathani et al. (2023); Gani et al. (2024).
Inspired by the self-refinement strategies by Madaan et al. (2023), we wanted to study whether the quality of our diagrams can be enhanced by using generic feedback for a wide range of errors or by applying targeted feedback for specific error categories. For exploring the second strategy, we introduce an iterative refinement framework, depicted in Figure 2 (Lower part), designed to enhance quality of diagram generation by specifically rectifying above-mentioned error types.
Overview of the Components:
As illustrated in Figure 3, our MAF framework comprises of the following components: a base model (SciDoc2Diagrammer) that generates an initial diagram ($O$) using both intent and source document (Step 1 in  3). It passes through three specialized feedback critic evaluator modules, each focusing on a critical aspect of diagram quality such as completeness ($C_{critic}$), faithfulness ($F_{critic}$), and layout ($L_{critic}$) on $D$, a Refiner module $R$ that generates refined diagram $O\textquoteright$ ’ based on $O$ and feedback.
Feedback Modules:
The completeness critic $C_{critic}$ (Algorithm 1) decomposes the diagram’s intent $I$ and generates a set of questions $Q$ from the user intent $I$ to determine how well various elements are represented in the diagram (Step 3). It then involves extracting answers from the diagram, scoring each response for adequacy, calculating an average score as the completeness score $C_{\text{Score}}$, and providing textual feedback ($C_{\text{Feedback}}$) to suggest improvements (Step 4).
The Faithfulness critic $F_{critic}$ (Algorithm 2) evaluates how accurately a diagram represents information from its document source. It evaluates the accuracy of a diagram by generating questions, retrieving answers from both the document and the diagram, and scoring their consistency to provide an average score and feedback for improvement. The layout critic $L_{critic}$ (Algorithm 4) evaluates how good is the look and feel of generated diagram, and provides score and feedback for improvement. We use two refinement techniques:
Summarization-Based Refinement Algorithm (SumMAF)
We adopt summarization-based multi-aspect refinement strategy from Nathani et al. (2023) for diagram refinement. However, in contrast to their method, we use our critic models that act as human proxies (Algorithm 4), and revisit PDF source and user intent repeatedly to validate and refine the diagram $C_{critic}$(Algorithm 1) and $F_{critic}$(Algorithm  2), mimicking the iterative human process of cross-verifying information through continuous questioning and answering.
The first draft is sent to $C_{critic}$ (Algorithm 1) which checks if all necessary components are present (In Figure 3, it checks the completeness of creating a flowchart). At the same time, the faithfulness evaluator $F_{critic}$ (Algorithm 2) ensures the accuracy of diagram by comparing it with source PDF, whether or not all the information in the source figure is true, by generating questions (FQ1, FQ2 in Figure 2) based on the diagram and obtaining answers from both PDF and diagram. Also, the layout evaluator checks the visual arrangement, ensuring good alignment and spacing (Figure 2). All the critic models provide feedback on a single output (Step 3’) and sent to the refiner. The refiner model makes adjustments by summarizing the suggested modifications from all critics and generates a new diagram (Step 4’). This is repeated (Steps 3’, 4’, 5’ in Figure 3) until the evaluator models are satisfied or we reach maximum iterations.
Sequential Refinement Algorithm (SeqMAF):
Algorithm 3 enhances a diagram in a step-by-step hierarchical manner ensuring that each critic is satisfied (See Figure 18). The intuition is that, complex tables or flowcharts benefit from being accurate and detailed in their depiction of processes or relationships, where sequential refinement allows for gradual addition of details, ensuring that each element is correctly represented. Here the diagram is feed into different critic modules one by one, refine the output until each of them are satisfied.
In Figure 3, $C_{critic}$ first provides feedback (detail composition operators or hierarchical representation, explicit modelling and long range dependency step inclusion) on the diagram in Step 4, and the refiner takes the initial figure and refines based on feedback in Step 5, and passed onto $C_{critic}$ to further evaluate and repeat the same steps (4, 5, 6) till maximum iteration is used or the $C_{critic}$ provides a score greater than 4.5 (as stated in Algorithm 1. The same process is repeated by $F_{critic}$ and $L_{critic}$ modules and generates output after Step 12.
5 Experimental Setup
Dataset Preprocessing.
Text on the papers in SciDoc2DiagramBench was extracted through scipdf Parser [^4] https://github.com/titipata/scipdf_parser which has been built on top of the Grobid GRO (2008–2024). Figures and captions were extracted through pdffigures2 Clark and Divvala (2016). We use 20 randomly selected diagrams from SciDoc2DiagramBench with stratified sampling to ensure that we have four type of diagrams (Table 1) for creating in-context examples in our prompts and the rest of the dataset is used for evaluation.
Flowchart Results Architecture Summary
Models (IV) R BS CS R BS CS R BS CS R BS CS
DALLE-3 0.12 0.45 0.23 0.11 0.34 0.21 0.25 0.45 0.25 0.04 0.10 0.18
Automatikz 0.20 0.48 0.25 0.20 0.50 0.29 0.24 0.45 0.28 0.27 0.56 0.45
SciDoc2Diagram
w/ GPT4-o (ZS) 0.22 0.58 0.43 0.32 0.49 0.30 0.24 0.50 0.34 0.33 0.61 0.62
w/ GPT4-o (FS) 0.28 0.67 0.43 0.40 0.56 0.38 0.32 0.57 0.45 0.45 0.67 0.62
w/ GPT4-o-SR 0.30 0.70 0.45 0.44 0.56 0.38 0.38 0.65 0.54 0.47 0.68 0.69
w/ GPT4-o-SumMAF 0.35 0.74 0.48 0.50 0.57 0.39 0.34 0.64 0.49 0.50 0.74 0.74
w/ GPT4-o-SeqMAF 0.39 0.79 0.53 0.49 0.49 0.36 0.37 0.58 0.49 0.45 0.67 0.62
Table 2: Automatic evaluation of models on various diagrams on SciDoc2DiagramBench-Gold using ROUGE (R), BERTScore (BS), CLIPScore (CS) with the highest values shown in green, highlighting that flowcharts, architectures are of the best quality after sequential refinement, whereas the other ones after summarization-based refinement.
Model Flowchart Summary Architecture Results
C F L C F L C F L C F L
GPT4-o w/o refinement 3.3 3.5 4.0 3.9 4.0 4.2 2.8 4.0 4.0 3.7 4.0 4.0
GPT4-o w/ refinement 3.9 4.0 4.0 4.5 4.1 4.2 3.2 4.0 4.0 3.7 4.5 3.8
Phi-3 w/o refinement 3.2 4.0 4.0 3.8 4.0 4.0 2.8 4.0 4.0 3.5 3.6 3.8
Phi-3 w/ refinement 3.8 4.0 4.3 4.4 4.1 4.2 3.0 4.3 4.0 3.8 4.1 3.8
Table 3: Comparison of Completeness (C), Faithfulness (F), and Layout (L) assessed by humans for different model variants on SciDoc2DiagramBench-Extended, indicating that on the complex set, Flowcharts and Summary tables improve significantly after refinement, but Layout-satisfaction is still comparatively lower.
Base LMs.
For generating first-draft of diagrams, we experiment with state-of-the-art models to extract table from existing figures during diagram planning (Figure 2): 1) MatCha Liu et al. (2023b), 2) Chart-to-Text Kantharaj et al. (2022), 3) UniChart Masry et al. (2023), 4) DePlot Liu et al. (2023a) and GPT-4 Vision (GPT-4V) OpenAI (2023). Finally, based on our assessment (Figure 20), we concluded that GPT-4V outperformed other models and was the best choice. For diagram planning, we experiment with three open-source LLMs: 1) LLama3 Huang et al. (2024), 2) Phi-3 Abdin et al. (2024) Mistral-7b Jiang et al. (2023) and for diagram rendering we use GPT4-o OpenAI (2023).
Critic Models for SciDoc2Diagrammer-MAF.
Inspired by the success of GPT-4V in making pairwise comparison of evaluation of images using Question-Answering driven approaches OpenAI (2023); Zhang et al. (2023) and image aesthetic evaluation Abe et al. (2024), we adopt GPT-4V model to provide feedback on different aspects of our generated diagrams on various criteria (Prompts can be found in 20,  21,  22).
Prior Work and Baselines.
We have compared our approaches with existing text-to-image generation methods like DALLE3 Rombach et al. (2021) and Automatikz Belouadi et al. (2023). For experimenting with Diagram Planning in SciDoc2Diagrammer, we evaluate using zero-shot and few-shot (with 3 in-context exemplars) versions of LMs as mentioned in Table 8 (Prompts in 15, 16, 18, 19). We use Self-Refine by Madaan et al. (2023) as another baseline (Prompt can be found in 25).
Figure 4: Average Human Rating on Completeness, Faithfulness and Layout on three subparts: SciDoc2DiagramBench-Gold (left), SciDoc2DiagramBench-Extended (middle) and SciMultiDoc2DiagramBench-Gold (right), it implies as the complexity of diagram creation increases, our proposed Multi-Aspect Refinement strategy appears to become more effective.
Evaluation Metrics.
SciDoc2DiagramBench-Gold and SciMultiDoc2DiagramBench-Gold include diagrams explicitly crafted by humans for each document or multiple documents, which can be served as a gold standard, allowing us to benchmark the quality of our generated diagrams. To evaluate our diagrams, we use three automated metrics: BERTScore Zhang et al. (2019), ROUGE-1 Lin (2004), and CLIPScore Hessel et al. (2021).
However, simple token evaluations and neural evaluations like BERTScore and CLIPScore are only measuring string similarity and lacking a more finegrained evaluation framework that correlates more with human preferences Li et al. (2024c). Thus, we go beyond string similarity scores and include a comprehensive analysis of more finegrained aspects through human evaluation and the GPT4-V Evaluation. We assess completeness (measures the degree to which all relevant and necessary information is in the generated figure), faithfulness (assesses how well the figure adheres to the facts, data, or specific instructions provided, ensuring that the content is correct and not misleading or misrepresented), and layout (measures visual clarity, focusing on how well the elements are structured and arranged) on a scale from 1 to 5.
Models (IV) FChart (R) FChart (BS) FChart (CS) Results (R) Results (BS) Results (CS) Arc (R) Arc (BS) Arc (CS) Summary (R) Summary (BS) Summary (CS)
w/ GPT4-o-SeqMAF 0.39 0.79 0.53 0.49 0.49 0.49 0.36 0.37 0.58 0.45 0.67 0.62
w/o Completeness-SeqMAF 0.25↓ 0.55↓ 0.35↓ 0.45 0.44 0.34↓ 0.36 0.57 0.47 0.30↓ 0.50↓ 0.45↓
w/o Faithfulness-SeqMAF 0.25↓ 0.55↓ 0.35↓ 0.35↓ 0.38↓ 0.25↓ 0.37 0.58 0.49 0.42 0.64 0.57
w/o Layout-SeqMAF 0.33 0.74 0.52 0.49 0.49 0.49 0.36 0.54 0.46 0.41 0.62 0.59
Table 4: Ablation Analysis Results: Here the column names FChart indicates the Extrapolated-Flowchart category of images, Arc indicates the Extrapolated-Architecture category of images, Results denotes the Extrapolated-Results category of images, and Summary denotes the Extrapolated-Summary category of images. Significant drop compared to the first row (greater than 0.1) is shown by ↓.
6 Results and Findings
Our main objective is to explore the creation of ‘Extrapolated’ scientific visuals from documents using zero-shot/few-shot settings of advanced VLMs, and also analyze whether the quality of these visuals without dataset-specific fine-tuning, can be improved using our proposed SciDoc2Diagrammer-MAF. To answer this broad question, we focus our experiments on answering the following research questions as follows:
6.1 Main Research Questions and Results
RQ1: What is the best base LM for generating diagrams before refinement?
GPT4-o is clearly the best-performing base/refiner model on human and automatic judgement. In Table 13, GPT-4o consistently outperforms other models across all categories of diagrams on the SciDoc2DiagramBench. This superior performance is indicated by the green cells, which show GPT-4o achieving the highest scores in all the automatic metrics. We observe similar trend when we ask annotators to rate on the completeness, faithfulness and layout-satisfaction on 30 randomly selected samples of images from SciDoc2DiagramBench-Gold generated by each model (Figure 19).
RQ2: How effective is our Scidoc2diagrammer-MAF over SciDoc2Diagrammer?
Based on Table 2, the effectiveness of our refinement strategies on SciDoc2DiagramBench-Gold using GPT4-o with SumMAF and SeqMAF enhancements shows significant improvements over the Zero-shot (ZS) and Few-shot (FS) versions on all the diagram types on automatic evaluation. This finding holds true for SciMultiDoc2DiagramBench-Gold as well, as evident from the results in Table 7.
Drawing inspiration from Ribeiro et al. (2020), we acknowledge that automatic metrics alone are insufficient to fully evaluate model performance, so we also conduct human evaluation. For SciDoc2DiagramBench Gold and SciMultiDoc2DiagramBench-Gold, we had human raters evaluate the quality of diagrams produced by different models and compare these to actual human-created slides, considering the source paper and the diagram’s intended purpose for 30 randomly sampled diagrams. In the SciMultiDoc2DiagramBench-Extended, which lacks a gold standard image, only the paper, the intent of the diagram, and the model-generated images were presented for evaluation on 50 randomly sampled images. Two authors and two professional crowdworkers from Upwork assessed the quality of all images, using criteria such as Completeness, Faithfulness, and Layout from 1 to 5. Figure 4 illustrates that diagrams generated using our refinement strategies compared to Few-shot and Self-Refine strategies consistently receive the highest scores from humans on all dataset subsets.
Figure 5: Refinement of a flowchart for detecting rumors from Ma et al. (2017), highlighting improvements in clarity and completeness after refinement but emphasizing the continued absence of a critical feedback loop.
RQ3: What type of refinement works the best across each diagram type based on automatic evaluation?
From Table 2, SeqMAF shows significant improvement with the highest scores in all metrics compared to both ZS and FS for flowcharts. SumMAF also improves over ZS and FS, but to a lesser extent than SeqMAF, indicating SeqMAF’s superior refinement capability on both completeness and detail accuracy. For Results diagrams, both SeqMAF and SumMAF show improvements over non-refined versions, with SeqMAF slightly leading in overall performance except for ClipScore. However, in Architecture and Summary diagrams, SumMAF occasionally outperforms SeqMAF, indicating its strength in contexts requiring synthesized summarization of content. Since summarization-based refinement offers an integrated approach where multiple facets of feedback are considered simultaneously, this might have been more effective for diagrams requiring balanced presentation of dense information. On SciDoc2DiagramBench-Extended, we explore whether human judgment perceives a significant enhancement in diagram quality across all categories on complex intents. We notice in Table 3, that even though there is a significant improvement of quality post-refinement on human judgement on Faithfulness and Completeness (Flowcharts and tables improve significantly), Layout-satisfaction is the least (which can be attributed to subjectivity and LLM’s inability to refine on specific criteria).
Figure 6: Qualitatively examines the images generated by different model-variants, where we show the incompleteness and non-faithfulness errors from our baseline models and how refinement improves the quality.
RQ4: Is refinement more effective on simple images or complex images?
Across three datasets, the Multi-Aspect Refinement strategy consistently outperforms the few-shot approach in diagram production (Figure 4). On the simplest dataset, SciDoc2DiagramBench-Gold, MAF-Refine performs significantly better than few-shot approach, with improvements of 0.9 in completeness, 1.1 in faithfulness, and 0.8 in layout. As complexity increases in SciDoc2DiagramBench-Extended, MAF-Refine still performs strongly, showing improvements of 1.4 in completeness, 1.0 in faithfulness, and 0.5 in layout. Even in the most complex dataset, SciMultiDoc2DiagramBench-Extended, MAF-Refine leads with significant improvements, showing that MAF-Refine significantly boosts diagram quality across datasets and maintains its benefits even as complexity rises.
6.2 Ablation Analysis
Impact of individual components in SciDoc2Diagrammer.
Table 9 presents the performance impact of different components in the SciDoc2Diagrammer model using different metrics. Removing the Question Generation/Answering (QG/QA) component results in significant performance drops across all metrics, highlighting its importance. The absence of the Summarization component also negatively affects the scores but to a lesser extent.
Impact of Refinements in SciDoc2Diagrammer.
We conduct ablation analysis (Table 4) with SciDoc2Diagrammer-SeqMAF. In particular, we experimented with the removal of each refinement module on the SciDoc2DiagramBench-Gold dataset, and analyzed which component is useful to improve the diagram generation. Completeness removal shows a significant drop in both summary and flowchart metrics, emphasizing the importance of comprehensive overviews and detailed visual representations. Faithfulness removal leads to a significant performance drop in results and flowcharts, underlining the necessity of faithful representation for coherent diagram creation. But layout removal has the least impact on diagram quality improvement.
6.3 Qualitative Analysis
Qualitative analysis in feedback-based refinement (Figure 5) identifies gaps in the initial design, such as missing components or unclear integration of elements. After refinement, while the layout and completeness may improve significantly, issues like the absence of critical feedback loops can persist, leading to a less effective refinement process. This shows that even with visible improvements, missing critical features can detract from the model’s intended dynamic refinement functionality.
Besides, taking a closer look at the diagrams generated before and after refinement (we sample the refined image having the highest average score rated by humans), we show in Figure  6,  21,  22 where our algorithm improves. For such qualitative analysis, we consider all four types of extrapolated images. The refined versions of the diagrams appear to show a higher level of detail and precision. For example, in the block diagrams and flowcharts, the GPT4-o with refinement diagrams show clearer layouts and better alignment of elements compared to those produced by baselines. The labels and connections between components are more accurate, suggesting that the refinement process improves the faithfulness and clarity of the diagrams. In scenarios where specific architectural details or complex data are necessary, the refined versions are closer to the original intent of diagram. This is particularly evident where GPT4-o with refinement accurately captures the necessary components and their interactions, unlike the outputs from baselines, which sometimes present a cluttered visualization.
7 Related Work
Our work connects to several distinct but interrelated fields, so we provide a comprehensive review of the most relevant prior work.
Text-to-Image Generation. Text-to-image generation models have seen immense success in recent years Ramesh et al. (2022b); Zala et al. (2023); Rombach et al. (2022b); Feng et al. (2023), but a handful of them support text conditioning Belouadi et al. (2023) on scientific images, and to the best of our knowledge, none of these explored challenges to incorporating in a long-text scenario.
Learning from Feedback. Madaan et al. (2023) first introduced ‘Self-Refine’ and then there had been multiple works exploring the use of natural language feedback to improve performance. Lee et al. (2024a) proposes a framework to reduce hallucination in question answering. Ki and Carpuat (2024) exploits the strengths of LLMs and machine translation by guiding LLMs to automatically post-edit MT with external feedback on its quality. Lee et al. (2024b) introduced a multimodal self-feedback guided revision model to reduce hallucination. The work most closely related to ours is that of Nathani et al. (2023). However, our approach diverges from theirs in a key aspect: our feedback modules are uniquely grounded in cross-verifying information from long texts. Furthermore, we introduce a novel algorithm that refines this process sequentially, which has proven to be more effective in our specific use case. We also discuss the related works on code generation in Appendix 18.
8 Discussion and Future Work
We introduce an innovative task for constructing scientific diagrams from scientific document discourse, commonly used in presentations/posters. We establish a new benchmark using document-to-slides dataset and a sequential diagram refinement-based methodology to generate diagrams from academic papers. In future, we would like to assess the granularity of feedback that helps in diagram improvement by designing rubrics about what a good diagram should look like by communicating with the experts in the data visualization communities.
Acknowledgement
We thank the anonymous ARR reviewers and UMD CLIP members for their constructive comments and feedback on the draft. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Annotations and support for Ishani Mondal were provided by a gift from Adobe Research.
Limitations
While our SciDoc2Diagrammer-MAF represents a significant advancement in automated diagram generation, some limitations should be noted:
  1. 1.
    Small Dataset Size. The task of creating diagrams that are both factually correct and visually appealing, based on long academic papers and user intent, requires deep domain knowledge and expert input. In the construction of SciDoc2DiagramBench, expert human annotators with extensive experience in natural language processing (NLP) and diagram creation were employed to ensure that the diagrams met the required standards of completeness, faithfulness, and layout quality. Hence, the current version of SciDoc2DiagramBench includes 1080 diagrams from 89 ACL papers, which, although diverse, is relatively small.
  2. 2.
    Complexity in Layout Refinement. As noted, our model struggles with improving the layout during the refinement process. This is particularly evident in complex visualizations where spatial and design elements are crucial. The current algorithms may not adequately capture aesthetic and functional aspects of layout design, impacting the overall clarity and effectiveness of the diagrams.
  3. 3.
    Feedback Mechanism Limitations. While the multi-aspect feedback mechanism aims to enhance diagram quality by focusing on completeness, faithfulness, and layout, the effectiveness of this feedback depends on the precision of the underlying evaluative criteria. Misalignments in feedback interpretation can lead to overfitting on specific diagram aspects at the expense of others.
Ethics Statement
The experiments performed in this study involved human participants. All the experiments involving human evaluation were exempt under institutional IRB review. We recruited participants for our human study using Upwork and we have fairly compensated all the Upwork freelancers involved in this study, at an average rate of $20.00$ USD per hour (respecting their suggested Upwork hourly wage). We did not collect any personal data during the experiments, and they could choose not to participate in the study. The documents used in the study are distributed under an open license.